HELP PLEASE!!!!!!!!!! use the substitution method......7x + 5y = -5
x= 37 - 7y

Answer by mananth(16946)

7x+5Y=-5
X=37-7Y
SUBSTITUTE X IN FIRST EQUATION
7(37-7y)+5y=-5
259-49y+5y=-5
-44y=-259-5
-44y=-264
/-44
y=6
plug y in second equation
x=37-7y
x=37-7*6
x=-5